# Non-Atomic Arbitrage in Decentralized Finance

### 15 Jan 2024 v2

@@ 6.4 Impact of Non-Atomic Arbitrage Trades
To conclude our analysis, we discuss the impact of non-atomic arbitrage and the vertical integration of searchers specializing in this arbitrage in the following. We start by noting that from 2023 onwards generally more than 50% of non-atomic arbitrage volume is found in HFT builder blocks (cf. Figure 11). Often this share rises to around 75%. Recall, that there are just four HFT builders, at no time were more than three of them in operation, and from May 2023 onwards only two were still operating (cf. Figure 14). Thus, we conclude that these builders have a grasp on the non-atomic arbitrage market.
@@
To further investigate the effects of cryptocurrency price volatility and thereby non-atomic arbitrage on the system, we take a closer look at the volatility in the lead-up to a block, i.e., for block $n$ the price volatility between the time the block $n-1$ and block $n$ were first seen in the network. On average (if blocks are not missed) this is twelve seconds on Ethereum. Increased volatility during this time should lead to increasingly profitable and prevalent non-atomic arbitrage transactions.
@@
Figure 13 plots the cdf of gas usage and block value of non-atomic arbitrage depending on the ETH and BTC price volatility. Observe that the higher the BTC and ETH price volatility in the lead-up to the block, the larger the share of the block’s gas usage by non-atomic arbitrage trades. Note that this is in addition to already larger blocks during times of high volatility (cf. Appendix B). Thus, the increased prevalence of non-atomic arbitrage likely leads to increased block competition during times of high volatility. Importantly, this also leads to an increased base fee (i.e., the minimum fee required for block inclusion) for consecutive blocks as the base fee automatically increases based on the previous block size. Further, this increase in the price for transaction inclusion as well as the overfullness of the block likely leads to increased congestion in the mempool. Thus, the block space used by non-atomic arbitrage during times of high volatility impacts all Ethereum users.
@@ 7 Discussion
In the following, we will highlight the implications of non-atomic arbitrage and hint at possible mitigation options.
7.1 Implication
@@
Previous works [6, 17] have demonstrated that MEV (i.e., high-value transactions) presents a risk to the consensus layer in PoW. To be exact, the consensus is vulnerable to time-bandit attacks, as it can be rational for the block proposer to fork the blockchain to exploit MEV in previous blocks themselves. Re-orgs, required by time-bandit attacks, have become harder in Ethereum PoS [29], but regardless such high-value transactions present a challenge to the consensus layers. For instance, an entity controlling a significant proportion of the staking power could purposefully withhold attestations blocks preceding its turn as block proposer to increase the chance of a possible re-org during its turn as proposer being successful. Note that the biggest staking pool currently controls around one-third of the staking power [30] and that the losses from missed attestations are minimal.
@@
Centralization of Block Building. We further comment on the centralizing effects of non-atomic arbitrage in the block construction market. Gupta et al. [4] have shown that HFT builders (i.e., builders with integrated non-atomic arbitrage searchers) are more likely to win blocks during times of high cryptocurrency for a one-month period in early 2023. We extend these findings to the entire history of Ethereum PoS until the end of October 2023 and highlight the relationships between builders and their integrated searchers. Our work further establishes a direct tie between the likelihood of these HFT builders winning the PBS auction and the volume of non-atomic arbitrage by their integrated searchers. As a result, during times of high cryptocurrency price volatility block building is largely left to two remaining HFT builders – a worrying centralization of the block construction market.
7.2 Mitigation
@@ A.1 Searcher - Builder Ties
TABLE IV: Sandwich attacked and cyclic arbitrage transactions included in blocks by the four builders identified to have integrated non-atomic arbitrage searchers.
To explain the lower correlation between the market shares of beaverbuild and its searchers in comparison to the other HFT builders, we take a look at the number of sandwich attacks and cyclic arbitrage transactions these builders include in their blocks. In Table IV find that beaverbuild and rsyncbuilder, where we also saw a lower correlation during the operation of rsyncseacher3, are much more likely to include sandwich attacks and cyclic arbitrage in their blocks, than builder1 and mantabuilder. Thus, we believe that the block value of beaverbuild and rsyncbuilder blocks also come from other types of MEV. Therefore, it is not only the non-atomic arbitrage volume that determines the value of their blocks even though it is a very significant driver.
@@ A.2 Beaverbuild
Appendix B Effects on Volatility
@@
Appendix C Heuristics
@@
To further ensure the validity of our heuristics, we investigate for each of the largest eleven searchers identified what proportion of their transactions we flag as non-atomic arbitrage in Table V. The previously identified non-atomic arbitrage searchers that we studied to develop these heuristics are included for completeness but marked in bold. We further note that these searchers are implemented as smart contracts and are likely to specialize in one profitable type of transaction. Thus, if we identify the majority of a searcher’s transactions as non-atomic arbitrage it is very likely that this classification is accurate. Turning to Table V, we notice that for all newly identified searcher except one (i.e., seacher2), more than 80% of their transactions fulfill all our five heuristics. This leads us to conclude that these searchers are in fact performing non-atomic arbitrage. We believe that we are more likely to undercount non-atomic arbitrage as the remaining swaps from these wallets are also likely to be (failed) non-atomic arbitrage swaps (i.e., they tend to be simple transfers and the first in the pool). However, we do not include these in our analysis to avoid overstating the volume of non-atomic arbitrage. Finally, we note that for searcher2 our heuristics only apply to 17% of swaps on DEXes. While the swaps we identify as non-atomic arbitrage appear to be exactly that this searcher also appears to perform other types of high-value transactions given that less than half of its swaps classify as simple.
Appendix D Builder and Seacher Addresses
We provide a mapping from searcher name to their address in Table VI, as well as mapping from builder names to their address(es) and public key(s) in Table VII.
name address
beaversearcher1 0xa69babef1ca67a37ffaf7a485dfff3382056e78c
beaversearcher2 0xa57bd00134b2850b2a1c55860c9e9ea100fdd6cf
builder1searcher 0x57c1e0c2adf6eecdb135bcf9ec5f23b319be2c94
jumpsearcher 0x9507c04b10486547584c37bcbd931b2a4fee9a41
rsyncsearcher1 0x0087bb802d9c0e343f00510000729031ce00bf27
rsyncsearcher2 0x280027dd00ee0050d3f9d168efd6b40090009246
rsyncsearcher3 0x51c72848c68a965f66fa7a88855f9f7784502a7f
mantasearcher 0xf8b721bff6bf7095a0e10791ce8f998baa254fd0
searcher1 0xe8cfad4c75a5e1caf939fd80afcf837dde340a69
searcher2 0x00000000008c4fb1c916e0c88fd4cc402d935e7d
searcher3 0xd7f3fbe8c72a961a5515203eada59750437fa762
searcher4 0x000000000dfde7deaf24138722987c9a6991e2d4
searcher5 0x98c3d3183c4b8a650614ad179a1a98be0a8d6b8e
TABLE VI: Searcher name and address mapping.
name address public key beaverbuild 0x95222290dd7278aa3ddd389cc1e1d165cc4bafe5 0x96a59d355b1f65e270b29981dd113625732539e955a1beeecbc471dd0196c4804574ff871d47ed34ff6d921061e9fc27 0xb5d883565500910f3f10f0a2e3a031139d972117a3b67da191ff93ba00ba26502d9b65385b5bca5e7c587273e40f2319 0x8dde59a0d40b9a77b901fc40bee1116acf643b2b60656ace951a5073fe317f57a086acf1eac7502ea32edcca1a900521 0xaec4ec48c2ec03c418c599622980184e926f0de3c9ceab15fc059d617fa0eafe7a0c62126a4657faf596a1b211eec347 0xacb407cfb554255db2fbbb320f79bb7f1cc1e8d2dc43324e8e31baafd0836340d49c43eebc51828f53bf6d364f9ac207 0xa21a2f4807a2bcb6b07c10cea241322e0910c30869c1e4eda686b0d69bdcb74d2a140ef994afcf0bb38e0b960df4d2ee bloXrouteM 0xf2f5c73fa04406b1995e397b55c24ab1f3ea726c 0x94aa4ee318f39b56547a253700917982f4b737a49fc3f99ce08fa715e488e673d88a60f7d2cf9145a05127f17dcb7c67 0x976e63c505050e25b70b39238990c78ddf0948685eb8c5687d17ba5089541f37dd3c45999f2db449eac298b1d4856013 0x8b8edce58fafe098763e4fabdeb318d347f9238845f22c507e813186ea7d44adecd3028f9288048f9ad3bc7c7c735fba 0xaa1488eae4b06a1fff840a2b6db167afc520758dc2c8af0dfb57037954df3431b747e2f900fe8805f05d635e9a29717b 0xafc9274fe595e8cff421ab9e73b031f0dff707ea1852e2233ff070ef18e3876e25c44a9831c4b5f802653d4678ccc31f bloXrouteR 0x199d5ed7f45f4ee35960cf22eade2076e95b253f 0x80c7311597316f871363f8395b6a8d056071d90d8eb27defd14759e8522786061b13728623452740ba05055f5ba9d3d5 0xb9b50821ec5f01bb19ec75e0f22264fa9369436544b65c7cf653109dd26ef1f65c4fcaf1b1bcd2a7278afc34455d3da6 0x965a05a1ba338f4bbbb97407d70659f4cea2146d83ac5da6c2f3de824713c927dcba706f35322d65764912e7756103e2 blocknative 0xbaf6dc2e647aeb6f510f9e318856a1bcd66c5e19 0x8000008a03ebae7d8ab2f66659bd719a698b2e74097d1e423df85e0d58571140527c15052a36c19878018aaebe8a6fea 0x9000009807ed12c1f08bf4e81c6da3ba8e3fc3d953898ce0102433094e5f22f21102ec057841fcb81978ed1ea0fa8246 0xa66f3abc04df65c16eb32151f2a92cb7921efdba4c25ab61b969a2af24b61508783ceb48175ef252ec9f82c6cdf8d8fd 0xa00000a975dffbd1ef61953ac6c90b52b70eb0188eb9d030774346c9248f81e875f7e8bc56c4bbbda297a9543cfa051d builder0x69 0x690b9a9e9aa1c9db991c7721a92d351db4fac990 0x8bc8d110f8b5207e7edc407e8fa033937ddfe8d2c6f18c12a6171400eb6e04d49238ba2b0a95e633d15558e6a706fbe4 0xb194b2b8ec91a71c18f8483825234679299d146495a08db3bf3fb955e1d85a5fca77e88de93a74f4e32320fc922d3027 0xa971c4ee4ac5d47e0fb9e16be05981bfe51458f14c06b7a020304099c23d2d9952d4254cc50f291c385d15e7cae0cf9d 0xa4fb63c2ceeee73d1f1711fadf1c5357ac98cecb999d053be613f469a48f7416999a4da35dd60a7824478661399e6772 0xb8fceec09779ff758918a849bfe8ab43cea79f6a98320af0af5b030f6a7850fcc5883cb965d02efb10eed1ffa987e899 builder1 0x473780deaf4a2ac070bbba936b0cdefe7f267dfc 0xa1daf0ab37a9a204bc5925717f78a795fa2812f8fba8bda10b1b27c554bd7dedd46775106facd72be748eea336f514e9 0x89783236c449f037b4ad7bae18cea35014187ec06e2daa016128e736739debeafc5fe8662a0613bc4ca528af5be83b3c flashbots 0xb64a30399f7F6b0C154c2E7Af0a3ec7B0A5b131a 0xdafea492d9c6733ae3d56b7ed1adb60692c98bc5 0x81babeec8c9f2bb9c329fd8a3b176032fe0ab5f3b92a3f44d4575a231c7bd9c31d10b6328ef68ed1e8c02a3dbc8e80f9 0x81beef03aafd3dd33ffd7deb337407142c80fea2690e5b3190cfc01bde5753f28982a7857c96172a75a234cb7bcb994f 0xa1dead01e65f0a0eee7b5170223f20c8f0cbf122eac3324d61afbdb33a8885ff8cab2ef514ac2c7698ae0d6289ef27fc mantabuiler 0x5f927395213ee6b95de97bddcb1b2b1c0f16844f 0xa0d0dbdf7b5eda08c921dee5da7c78c34c9685db3e39e81eb91da94af29eaa50f1468813c86503bf41b4b51bf772800e 0xb1b734b8dd42b4744dc98ea330c3d9da64b7afc050afed96875593c73937d530a773e35ddc4b480f9d2e1d5ba452a469 0xb5a688d26d7858b38c44f44568d68fb94f112fc834cd225d32dc52f0277c2007babc861f6f157a6fc6c1dc25bf409046 rsyncbuilder 0x1f9090aae28b8a3dceadf281b0f12828e676c326 0x978a35c39c41aadbe35ea29712bccffb117cc6ebcad4d86ea463d712af1dc80131d0c650dc29ba29ef27c881f43bd587 0x83d3495a2951065cf19c4d282afca0a635a39f6504bd76282ed0138fe28680ec60fa3fd149e6d27a94a7d90e7b1fb640 0xacfdcf458829f4693168a57d0659253069d687682bc64ec130d935ecb6e05ccfb80c138bd3cf53546c86715696612ec8 0x945fc51bf63613257792926c9155d7ae32db73155dc13bdfe61cd476f1fd2297b66601e8721b723cef11e4e6682e9d87 0x8aab0ed724d2c7f94af139bd2249ab511f08474ac69e761e56918403c81c358f5f8a6d61c62a86dc4cd7bcad935f49d9 0x8e6df6e0a9ca3fd89db2aa2f3daf77722dc4fbcd15e285ed7d9560fdf07b7d69ba504add4cc12ac999b8094ff30ed06c
TABLE VII: Builder name, address(es), and public key(s).
Appendix E Cryptocurrencies
We provide a list of the 60 cryptocurrencies tracked in this study in Table VIII chosen based on most interactions on DEXes as well as price information on CEXes.
name address
USDC 0xa0b86991c6218b36c1d19d4a2e9eb0ce3606eb48
USDT 0xdac17f958d2ee523a2206206994597c13d831ec7
(W)ETH 0xc02aaa39b223fe8d0a0e5c4f27ead9083c756cc2
DAI 0x6b175474e89094c44da98b954eedeac495271d0f
(W)BTC 0x2260FAC5E5542a773Aa44fBCfeDf7C193bc2C599
LDO 0x5a98fcbea516cf06857215779fd812ca3bef1b32
LINK 0x514910771af9ca656af840dff83e8264ecf986ca
UNI 0x1f9840a85d5af5bf1d1762f925bdaddc4201f984
PEPE 0x6982508145454ce325ddbe47a25d4ec3d2311933
CRV 0xd533a949740bb3306d119cc777fa900ba034cd52
APE 0x4d224452801aced8b2f0aebe155379bb5d594381
AAVE 0x7fc66500c84a76ad7e9c93437bfc5ac33e2ddae9
MATIC 0x7d1afa7b718fb893db30a3abc0cfc608aacfebb0
MKR 0x9f8f72aa9304c8b593d555f12ef6589cc3a579a2
SHIB 0x95ad61b0a150d79219dcf64e1e6cc01f0b64c4ce
QNT 0x4a220e6096b25eadb88358cb44068a3248254675
RPL 0xd33526068d116ce69f19a9ee46f0bd304f21a51f
COMP 0xc00e94cb662c3520282e6f5717214004a7f26888
ENS 0xc18360217d8f7ab5e7c516566761ea12ce7f9d72
SUSHI 0x6b3595068778dd592e39a122f4f5a5cf09c90fe2
SYN 0x0f2d719407fdbeff09d87557abb7232601fd9f29
LQTY 0x6dea81c8171d0ba574754ef6f8b412f2ed88c54d
ILV 0x767fe9edc9e0df98e07454847909b5e959d7ca0e
SNX 0xc011a73ee8576fb46f5e1c5751ca3b9fe0af2a6f
ALCX 0xdbdb4d16eda451d0503b854cf79d55697f90c8df
CVX 0x4e3fbd56cd56c3e72c1403e103b45db9da5b9d2b
dYdX 0x92d6c1e31e14520e676a687f0a93788b716beff5
1INCH 0x111111111117dc0aa78b770fa6a738034120c302
AGIX 0x5b7533812759b45c2b44c19e320ba2cd2681b542
FXS 0x3432b6a60d23ca0dfca7761b7ab56459d9c964d0
SPELL 0x090185f2135308bad17527004364ebcc2d37e5f6
YFI 0x0bc529c00c6401aef6d220be8c6ea1667f6ad93e
BAT 0x0d8775f648430679a709e98d2b0cb6250d2887ef
GALA 0x15d4c048f83bd7e37d49ea4c83a07267ec4203da
WLD 0x163f8c2467924be0ae7b5347228cabf260318753
CREAM 0x2ba592f78db6436527729929aaf6c908497cb200
REN 0x408e41876cccdc0f92210600ef50372656052a38
FTM 0x4e15361fd6b4bb609fa63c81a2be19d873717870
RNDR 0x6de037ef9ad2725eb40118bb1702ebb27e4aeb24
OGN 0x8207c1ffc5b6804f6024322ccf34f29c3541ae26
NEXO 0xb62132e35a6c13ee1ee0f84dc5d40bad8d815206
PERP 0xbc396689893d065f41bc2c6ecbee5e0085233447
IMX 0xf57e7e7c23978c3caec3c3548e3d615c346e79ff
MANA 0x0f5d2fb29fb7d3cfee444a200298f468908cc942
KP3R 0x1ceb5cb57c4d4e2b2433641b95dd330a33185a44
YGG 0x25f8087ead173b73d6e8b84329989a8eea16cf73
ID 0x2dff88a56767223a5529ea5960da7a3f5f766406
SAND 0x3845badade8e6dff049820680d1f14bd3903a5d0
BLUR 0x5283d291dbcf85356a21ba090e6db59121208b44
OCEAN 0x967da4048cd07ab37855c090aaf366e4ce1b9f48
stETH 0xae7ab96520de3a18e5e111b5eaab095312d7fe84
FET 0xaea46a60368a7bd060eec7df8cba43b7ef41ad85
PRIME 0xb23d80f5fefcddaa212212f028021b41ded428cf
HFT 0xb3999f658c0391d94a37f7ff328f3fec942bcadc
BAND 0xba11d00c5f74255f56a5e366f4f77f5a186d7f55
GRT 0xc944e90c64b2c07662a292be6244bdf05cda44a7
GALA 0xd1d2eb1b1e90b638588728b4130137d262c87cae
ZRX 0xe41d2489571d322189246dafa5ebde1f4699f498
GALA 0xd1d2eb1b1e90b638588728b4130137d262c87cae
ENJ 0xf629cbd94d3791c9250152bd8dfbdf380e2a3b9c
HEX 0x2b591e99afe9f32eaa6214f7b7629768c40eeb39
TABLE VIII: List of cryptocurrencies tracked.

### 08 Apr 2024 v3

@@ 6.4 Impact of Non-Atomic Arbitrage Trades
To conclude our analysis, we discuss the impact of non-atomic arbitrage and the vertical integration of searchers specializing in this arbitrage. We start by noting that from 2023 onwards generally more than 50% of non-atomic arbitrage volume is found in HFT builder blocks (cf. Figure 11). Often this share rises to around 75%. Recall, that there are just four HFT builders, at no time were more than three of them in operation, and from May 2023 onwards only two were still operating (cf. Figure 14). Thus, we conclude that these builders have a strong grasp on the non-atomic arbitrage market.
@@
To further investigate the effects of cryptocurrency price volatility and thereby non-atomic arbitrage on the system, we take a closer look at the volatility in the lead-up to a block, i.e., for block $n$ the price volatility between the time block $n-1$ and block $n$ were first seen in the network. On average (if blocks are not missed) this is twelve seconds on Ethereum. Increased volatility during this time should lead to increasingly profitable and prevalent non-atomic arbitrage transactions.
@@
Figure 13 plots the cdf of gas usage and block value of non-atomic arbitrage depending on the ETH and BTC price volatility. Observe that the higher the BTC and ETH price volatility in the lead-up to the block, the larger the share of the block’s gas usage by non-atomic arbitrage trades (cf. Figure 12(a)). Overall, in less than 1% of blocks, non-atomic arbitrage transactions make up more than 5% of the block’s total gas usage (red line). However, for blocks above the 99.9th percentile of BTC price volatility (yellow line), we see that in around 40% of blocks, non-atomic arbitrage transactions use more than 10% of block space, i.e., gas. Note that this is in addition to already larger blocks during times of high volatility (cf. Appendix B). Thus, the increased prevalence of non-atomic arbitrage likely leads to increased block competition during times of high volatility. Importantly, this also leads to an increased base fee (i.e., the minimum fee required for block inclusion) for consecutive blocks as the base fee automatically increases based on the previous block size. Further, this increase in the price for transaction inclusion as well as the overfullness of the block likely leads to increased congestion in the mempool. Thus, the block space used by non-atomic arbitrage during times of high volatility impacts all Ethereum users.
@@ 7 Discussion
Previous works [6, 18] have demonstrated that MEV (i.e., high-value transactions) presents a risk to the consensus layer in PoW. To be exact, the consensus is vulnerable to time-bandit attacks, as it can be rational for the block proposer to fork the blockchain to exploit MEV in previous blocks themselves. Re-orgs, required by time-bandit attacks, have become harder in Ethereum PoS [30], but regardless such high-value transactions present a challenge to the consensus layers. For instance, an entity controlling a significant proportion of the staking power could purposefully withhold attestations for blocks preceding its turn as block proposer to increase the chance of a possible re-org during its turn as proposer being successful. Note that the biggest staking pool currently controls around one-third of the staking power [31] and that the losses from missed attestations are minimal.
@@
Centralization of Block Building. We further comment on the centralizing effects of non-atomic arbitrage in the block construction market. Gupta et al. [4] have shown that HFT builders (i.e., builders with integrated non-atomic arbitrage searchers) are more likely to win blocks during times of high cryptocurrency price volatility for a one-month period in early 2023. We extend these findings to the entire history of Ethereum PoS until the end of October 2023 and highlight the relationships between builders and their integrated searchers. Our work further establishes a direct tie between the likelihood of these HFT builders winning the PBS auction and the volume of non-atomic arbitrage by their integrated searchers. As a result, during times of high cryptocurrency price volatility block building is largely left to two remaining HFT builders – a worrying centralization of the block construction market.
8 Mitigation
@@ A.1 Searcher - Builder Ties
TABLE IV: Sandwich attack and cyclic arbitrage transactions included in blocks by the four builders identified to have integrated non-atomic arbitrage searchers.
To explain the lower correlation between the market shares of beaverbuild and its searchers in comparison to the other HFT builders, we take a look at the number of sandwich attacks and cyclic arbitrage transactions these builders include in their blocks. In Table IV find that beaverbuild and rsyncbuilder, where we also saw a lower correlation during the operation of rsyncseacher3, are much more likely to include sandwich attacks and cyclic arbitrage in their blocks, than builder1 and mantabuilder. Thus, we believe that the block value of beaverbuild and rsyncbuilder blocks also come from other types of MEV. Therefore, it is not only the non-atomic arbitrage volume that determines the value of their blocks even though there is a significant correlation.
@@ A.2 Beaverbuild
Appendix B Effects of Volatility
@@
Appendix C MEV
(a) Daily number of occurrences of non-atomic arbitrage in comparison to sandwich attacks, cyclic arbitrage and liquidations.
(b) Daily value of of non-atomic arbitrage in comparison to sandwich attacks, cyclic arbitrage and liquidations.
Figure 17: Comparison between non-atomic arbitrage and well-studied types of MEV, i.e., sandwich attacks, cyclic arbitrage, and liquidations.
In Figure 17, we provide a comparison between non-atomic arbitrage and well-studied types of MEV, i.e., sandwich attacks, cyclic arbitrage, and liquidations. First, we compare the daily number of occurrences of each type of MEV in Figure 16(a). Note that for sandwich attacks, which consist of a frontrunning transaction, the victim’s transaction, and a backrunning transaction, we count this set of transactions as a single transaction. Non-atomic arbitrage and sandwich attacks dominate MEV in terms of a number of occurrences. Additionally, in the first part of our data collection window, the number of occurrences of non-atomic arbitrage exceeds that of sandwich attacks. Then around May 2023, it switches and there are more sandwich attacks than occurrences of non-atomic arbitrage. Notably, this coincides with jaredfromsubway.eth, i.e., a sandwich attack bot, entering the MEV market. By July 2023, the occurrences of non-atomic arbitrage and sandwich attacks are quite similar and continue to be so until the end of our data collection window.
We further compare the value, i.e., the fees paid to the proposer/builder through priority fees and coinbase transfers, of the four types of MEV (cf. Figure 16(b)). Again we notice that non-atomic arbitrage and sandwich attacks are responsible for the majority of the value. To put it in numbers, the average daily non-atomic arbitrage value is 112 ETH, whereas it is 161 ETH for sandwich attacks, 26 ETH for cyclic arbitrage, and 5 ETH for liquidations. We again see an increase in the value of sandwich attacks, as jaredfromsubway.eth enters the MEV market around May 2023, which then declines again. Additionally, we observe that on days with high cryptocurrency price volatility, such as the FTX collapse in November 2022 or the USDC depeg in March 2023, we observe spikes in all types of MEV.
To conclude, the scale of non-atomic arbitrage in terms of number of occurrences and in terms of value is comparable to that of sandwich attacks and significantly exceeds that of cyclic arbitrage and liquidations.
Appendix D Heuristics
@@
To further ensure the validity of our heuristics, we investigate the portion of transactions we flag as non-atomic arbitrages in Table V for each of the largest eleven searchers that we identified. The previously identified non-atomic arbitrage searchers that we studied to develop these heuristics are included for completeness and marked in bold. These searchers were chosen as Gupta et al. [4] previously showed that these searchers were likely involved in CEX-DEX arbitrages. We further note that these searchers are implemented as smart contracts and are likely to specialize in one profitable type of transaction. Thus, if we identify the majority of a searcher’s transactions as non-atomic arbitrage it is very likely that this classification is accurate. Turning to Table V, we notice that for all newly identified searcher except one (i.e., seacher2), more than 80% of their transactions fulfill all our five heuristics. This leads us to conclude that these searchers are in fact performing non-atomic arbitrage. We believe that we are more likely to undercount non-atomic arbitrage as the remaining swaps from these wallets are also likely to be (failed) non-atomic arbitrage swaps (i.e., they tend to be simple transfers and the first in the pool). However, we do not include these in our analysis to avoid overstating the volume of non-atomic arbitrage. Finally, we note that for searcher2 our heuristics only apply to 17% of swaps on DEXes. While the swaps we identify as non-atomic arbitrage appear to be simple swaps, this searcher also appears to perform other types of high-value transactions given that less than half of its swaps classify as simple.